NVP versus IRR. Romboski, LLC, has identified the following two mutually exclusive projects:

Year

Cash Flow (A)

Cash Flow (B)

0

($65,000)

($65,000)

1

$34,000

$19,000

2

$27,000

$25,000

3

$21,000

$29,000

4

$17,000

$34,000

IRR

22.23%

21.01%

  1. Over what range of discount rates would you choose Project A? Project B? At what discount rate would you be indifferent between these two projects? Explain.
Answer #1

fi =IRR(B21:B25) B26 А 1 A 2 year в Cash flows pv@16.31% Present value 0 $(65,000.00) 1.0000 $ (65,000.00) 1 $ 34,000.00 0.85B26 fx =IRR(B21:B25) А pv@16.31% 1 A 2 year 30 4 1 5 2 63 74 8 IRR Cash flows -65000 34000 27000 21000 17000 =IRR(B3:37) =C3/